# Probabilistic Counters for Privacy Preserving Data Aggregation11footnotemark: 1

### v5

@@ 8.1 Main Theorem for Morris Counter (Theorem 1)
Actually, we may consider ${}^{\prime}+^{\prime}$ sign instead of ${}^{\prime}\pm^{\prime}$ in (10), because $|\ln(x)|=|\ln({1}/{x})|$. However, when $I_{n}\neq I_{n\pm 1}$, we have to behave carefully, so in particular, an additional check of privacy loss with ${}^{\prime}-^{\prime}$ sign is needed when $n$ is of a form $12^{l}+$ 1 for some $l\in\mathbb{N}$.
@@
This claim emerges from lemmas 7 and 8 from A.
@@
We use Claim 2 to guarantee starting conditions for the next Lemma 3. However, in order to apply Lemma 3, we will use Claim 1 as well, which assumes that $n\geqslant 2^{7}+1$. Hence we would like to gather some information about the distribution of $M_{2^{7}+1}$. More precisely, we are interested in the behaviour of $\theta_{i}={p_{129,i}}/{p_{129,i+1}}$ for $i\leqslant 11$, which we present in the Table 1.
@@
We briefly see a superexponential trend of proportions $\theta_{i}$, so a possibility to use Claim 2 for $n\geqslant 2^{7}+1$ is justified. It might seem that the choice of $n$ is arbitrary, but it occurs that the distribution of $M_{2^{6}+1}$ does not have satisfy the necessary assumptions for the privacy loss although $M_{2^{6}+1}$ still can fulfill the property of $(\varepsilon(n),\delta)$-DP with parameters given in Theorem 1.
@@
Let $k\in\mathbb{N}\setminus\{0\}$ and $n_{k}=2^{k}+1$. If $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$ for every $c$ in the interval $[-k:4]$, then
@@
Realize that for $c=-k$, the required inequality is trivial. Therefore we can safely consider only $c\in[-k+1:4]$. We would like to prove it inductively with respect to $c$ and $N$. Assume that for some $N\geqslant n_{k}$ and any $d\in\{c-1,c\}$ we have $p_{N,k+d}<2^{d+3}p_{N,k+d+1}$. Then also
@@
If we start with $c=-k+1$, then the above let us prove inductively the appropriate condition for all $N\geqslant n_{k}$. Further the thesis is followed by the induction with respect to $c$$\qed$
@@
According to the previous discussion about the formula (10), we examine:
@@
Realize that if $n=2^{l-1}+1$ for some $l\in\mathbb{N}$, then a little adjustment is necessary. Indeed, let now $c-1=k-l\in[-4:4]$, and once again, Lemma 3 provides $p_{n-1,k-1}<~{}2^{c+2}p_{n-1,k}$. However, it still holds that:
@@
On the other hand, we obviously have inequalities $p_{n+1,k}>\left(1-2^{-l-c}\right)p_{n,k}$ and $p_{n,k}>\left(1-2^{-l-c}\right)p_{n-1,k}$ for any $c\in[-4:4]$, so both of these fractions exceed $1-{16}/{n}$$\qed$
Theorem 6 only provides $\varepsilon(n)\leqslant-\ln\left(1-{16}/{n}\right)$ for $n>128$ (compare with (10)). However, in the Figure 5 we may briefly see that the above inequality is true for smaller number of requests $n$ as well.
@@
hence from Fact 1 we obtain the main result.$\qed$
Figure 5: Exact values of $\varepsilon(n)$ parameter for $n\leqslant 160$ compared with plots of sequences $-\ln(1-{16}/{n})$ and $-\ln(1-{8}/{n})$.
In the Figure 5 one can see that values of $\varepsilon(n)$ are strictly between sequences $-\ln(1-~{}{8}/{n})$ and $-\ln(1-~{}{16}/{n})$ for $n\in[17:160]$. We can also observe that $\varepsilon(n)~{}\approx~{}2^{4-\lceil\log(n)\rceil}$ in this interval. Note that $\lceil\log(n)\rceil\leqslant 4$ for $n\leqslant 16$, so $\lceil\log(n)\rceil-4<1$, but $M$ is always positive. This can justify the chaotic behavior of the process for $n\leqslant 16$. Nevertheless, Figure 5 affirms the quality of $\varepsilon(n)$ parameter established in Theorem 1. Moreover, let us mention, that for the strategy used in the proof, we cannot pick smaller $\varepsilon$ of the same form:
@@
As may be seen in Figure 5, minimal $c$ for $\varepsilon(n)$, where $n$ is of the form $2^{k}$, for $k>5$, is very close to $16$ as well.
@@ 8.2 General result on Morris’ Counter privacy (Theorem 2)
For our convenience, let us denote $\rho:=\lceil c\log(\ln(n))\rceil$. We assume that $\rho\geqslant 1$. First we show that $\delta_{1}^{\ast}:=\mathbb{P}\left(M_{n}\leqslant\lceil\log(n)\rceil-\rho-1\right)=O\left(n^{-\left(\ln(n)\right)^{c-1}}\right)$. The proof is analogous to the one of Lemma 1 (we omit the similar parts). Indeed,
@@
Now, we are going to prove $\delta_{2}^{\ast}:=\mathbb{P}\left(M_{n}\geqslant\lceil\lg(n)\rceil+\rho+1\right)=O\left(n^{-1}\left(\ln(n)\right)^{-c}\right)$. We use similar notation and technique as in proof of Lemma 2, but this time we utilize $(\rho-1)$-dimensional discrete simplex:
@@
We couple $(M_{n})_{n}$ with the same process $(X_{n})_{n}$ as in Lemma 2. Roughly speaking, $X_{0}=1$ and $X_{n}$ almost always increments by $1$ for until $n=\lceil\log(n)\rceil+1$ and further it follows the same rule of incrementation as Morris counter. Then
@@
Further we would like to consider fractions $\frac{p_{n+1,k}}{p_{n,k}}$ for $k\in J_{n}(c)$ as in the proof of Theorem 6. Indeed
@@
Let us denote the above sum by $\varsigma_{k}(n-k+1)$. Let us note that
@@
Realize that $\varsigma_{k-1}(n-k+2)\leqslant\varsigma_{k-1}(n-k+1)\sum_{i=1}^{k-1}\left(1-2^{-i}\right)=\varsigma_{k-1}(n-k+1)\left(k-2+2^{k-1}\right)$. This follows from the fact that each summand of $\varsigma_{k-1}(n-k+2)$ can be obtained from some summands of $\varsigma_{k-1}(n-k+1)$ by multiplication by one of the terms $\left(1-2^{-i}\right)$. Moreover, note that $\varsigma_{k}(n-k+1)$ has $\binom{n-k+1+(k-1)}{k-1}=\binom{n}{k-1}$ summands and similarly, $\varsigma_{k-1}(n-k+1)$ has $\binom{n-1}{k-2}$ summands. One can briefly see that a function $f(i)=\left(1-2^{-i}\right)$ is increasing, hence
@@
Due to the monotonicity of $f$, one can use cascading substitutions: some of $f(k)$ by $f(k-1)$, then some of $f(k-1)$ by $f(k-2)$ etc., in order to balance the numbers of all the occurring summands, what provides:
@@ 8.3 Main result for MaxGeo Counter (Theorem 3)
We have $n$ increment requests, which influence the value of MaxGeo Counter $M$. Thence the result of the mechanism is $X=\max(X_{1},\ldots,X_{n})$, where $X_{i}\sim\mathrm{Geo}({1}/{2})$ are pairwise independent. First, we observe that if $n=m$, the counter trivially satisfies differential privacy, as the probability distribution of $X$ does not change. From now on, we assume that $|n-m|=1$. See that
@@
It is easy to see that the above decreases when $n$ increases. Then
@@
where the approximation is a result of substitution of $l_{\varepsilon}$ without ceiling. $\qed$
@@ 9 Conclusions and Future Work
In this paper, we have investigated probabilistic counters from the privacy-protection perspective. We have shown that Morris Counter and MaxGeo Counter inherently guarantee differential privacy from the mechanism itself, provided that there is at least a small, fixed number of increment requests. Otherwise, the counter has too low value, and intuitively, the result is not randomized enough. We have also shown that the constant in our Morris Counter result cannot be improved further.
@@
We have shown how to perform data aggregation, namely a distributed survey, in a privacy-preserving manner using probabilistic counters. We clarified that this type of solution is especially efficient when one cares about memory resources, like in many Big Data related problems. Note that the security model in this paper was somewhat optimistic. Unfortunately, in such a setting, there is little incentive to use them other than when we already have them deployed and working as aggregators due to e.g., memory-efficiency requirements. This would, however, change tremendously if we would weakened these assumptions. This seems a very promising way to continue our research from this paper. Namely, we focused on privacy and can still not weaken the security assumptions and allow the Adversary to extract information from channels between users and the aggregator. That would put us in the so-called, Local Model, where each user is responsible for the data randomization. However, such an approach requires us to be able to perform probabilistic counter in an oblivious manner which, to the best of our knowledge, was not explored before.
@@
In Subsection 4.1, we have mentioned the generalization of Morris Counter (for bases $a>1$). Analysis of privacy properties of such variants of Morris Counters and various probabilistic counters presented, for example, in [28], [48] may also be promising direction of further research.
@@
In this paper, we focused on the standard definition of differential privacy. However, there is also an issue of preservation of differential privacy for requests given by a group of $k$ users or one individual sending up to $k$ dependent requests over time. This can be described in a language of so-called $(\varepsilon,\delta)-k$-DP (see "group privacy scenario" in [5]) . A group of people may tend to behave in the same manner, so they may send $k$ requests in a row. Especially this ”group” may be represented by a single person colluding with the Adversary. It is worth mentioning that this type of generalization creates an opportunity to modify probabilistic counters so that each incrementation request executes the update request multiple times to reduce the variance of the rescaled estimator. Intuitively, this extension should be especially efficient in preserving standard differential privacy property when $\varepsilon(n)=\frac{c}{n}+o(n^{-1})$ (as a parameter of standard differential privacy), because both $c$ and $n$ should scale with $k$ linearly. Hence the next challenging problem is showing that Morris and MaxGeo Counters satisfy $k$-DP property with similar privacy parameters.
@@
Morris Counter and MaxGeo Counter are considered the most popular probabilistic counters. However, the results of this paper shed new light on the properties of the probabilistic counter in general. There is a possibility to provide analogous differential privacy properties for other probabilistic counters. Moreover, this paper enables the provision of differentially private algorithms for other applications, especially those based on Morris or MaxGeo Counter. For instance, in Section 4 we mentioned PCSA and HyperLogLog Counters together with their variances, which can be manually adjusted to applications. A proper choice of $m$ parameter implies an exchange of the memory usage for the accuracy of the estimation. We have mentioned that these counters’ differential privacy parameters can be obtained via Observation 1. However, such a direct result may not be satisfying. Hence a more precise calculation is needed. For instance, Observation 1 may be utilized again with some concentration inequalities.
@@ Acknowledgement
This work was supported by Polish National Science Center grant number UMO-2018/29/B/ST6/02969.
@@ Appendix A Technical Lemmas and Proofs Related to Differential Privacy of Morris Counter
For the sake of completeness, we present here proofs of all technical lemmas that are not directly connected to Theorem 1. Some of computations are supported by Wolfram Mathematica ver.11.3 ([49]). Whenever we obtain a result in this manner, we indicate it by =$\stackrel{{\scriptstyle W}}{{=}}$ sign. Usually results are precise, however in some cases, final forms are attained numerically. We often struggle with expressions of a pattern $1-{1}/{y}$, so let us denote this function as $a(y)$ to abbreviate formulas.
@@
and in a consequence $a(2^{-s}x)^{2x+1}<\exp(-2^{s+1})$ for any reasonable $s$. Moreover, let us introduce
@@
Let $d:=1-{2^{s}}/{x}$ and realize that $d\in[{3}/{4},1)$ and $02^{s}-1+d\left(-{2^{s}}/{3}-2^{-6}-2^{2s-2}\right)>$ 0. Indeed, if we put $z=2^{s}$, then we attain a quadratic inequality in $z$ variable, with determinant $\Delta=1-\frac{5d}{3}+\frac{55d^{2}}{576}$, that is negative for $d\in[{3}/{4},1)$.
@@
$\tau(k,t):=[\![2\!\!\not|(k+t)]\!](-1)^{k+t+3}2^{-\frac{(k+t+3)(k+t+2)}{2}}2r_{k+t+3}\left(\left(\frac{3}{4}\right)^{2x+1}-\left(\frac{1}{2}\right)^{x+2}\right)~{},$
where $[\![\mathrm{cond}]\!]$ is the Iverson bracket of the condition $\mathrm{cond}$. Realize that for $t\in\{0,1\}$ and $k\geqslant 5$, $|\tau(k,t)+\kappa(k,t)|<2^{-50}<10^{-15}$. Now, consider the differences between the consecutive elements of sequences:
@@
$=$ $\sum_{i=0}^{\left\lfloor\frac{k+t+2}{2}\right\rfloor}2^{-i(2i-1)}r_{2i+1}r_{k+t+4-2i}\left[a(2^{2i+1})\left(a(2^{5+t-2i}x)^{2x+1}-a(2^{4+t-2i}x)^{x+2}\right)\right.$
@@
=$\stackrel{{\scriptstyle W}}{{=}}a(2^{2i+1})a(u_{t}x)^{2x+1}\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{u_{t}+1}{u_{t}^{3}x}-\frac{1}{u_{t}^{3}x^{2}}\right)$
@@
=$\stackrel{{\scriptstyle W}}{{=}}a(2^{2i+1})\exp\left(-\frac{2}{u_{t}}\right)\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{u_{t}+1}{xu_{t}^{3}}-\frac{1}{x^{2}u_{t}^{3}}\right)$ (12)
@@
=$\stackrel{{\scriptstyle W}}{{=}}$ $a(2^{2i+1})a\left(\frac{u_{t}}{2}x\right)^{x-1}\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{10u_{t}+1}{u_{t}^{3}x}+\frac{8}{u_{t}^{3}x^{2}}\right)$ (13)
@@
=$\stackrel{{\scriptstyle W}}{{=}}\frac{a(2^{2i+1})}{x}\exp\left(-\frac{2}{u_{t}}\right)\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{2+5u_{t}+4u_{t}^{2}}{u_{t}^{4}x}-\frac{2+4u_{t}-\frac{575}{32}u_{t}^{2}+\frac{387}{32}u_{t}^{3}}{u_{t}^{6}x^{2}}\right.$
@@
(16$)\geqslant\frac{C(t,x)}{256}\left(31(14u_{t}+7)-96u_{t}-64\right)>0$. When $i=1$, then $u_{t}\geqslant 8$, so (16$)\geqslant\frac{C(t,x)}{64}\left(7(14u_{t}+7)-96u_{t}-64\right)=\frac{2u_{t}-15}{64}\geqslant\frac{1}{64}$.
@@
[…]
@@
However, according to Theorem 4, we also present the numerical values of the sequence $(p_{2^{k}+1,k+4})_{k=2}^{14}$ in the Table 2. We can now easily see that for any $k\geqslant 2$ we attained $p_{2^{k+1}+1,k+5}-p_{2^{k}+1,k+4}<0$

### v6

@@ 8.1 Main Theorem for Morris Counter (Theorem 1)
Actually, we may consider the sign $^{\prime}+^{\prime}$ instead of $^{\prime}\pm^{\prime}$ in (10), because $|\ln(x)|=|\ln({1}/{x})|$. However, when $I_{n}\neq I_{n\pm 1}$, we have to behave carefully, so in particular, an additional cheque of privacy loss with the sign $^{\prime}-^{\prime}$ is needed when $n$ is of a form $2^{l}+1$ for some $l\in\mathbb{N}$.
@@
This claim arises from lemmas 7 and 8 from A.
@@
We use Claim 2 to guarantee starting conditions for the next Lemma 3. However, in order to apply Lemma 3, we will also use Claim 1, which assumes that $n\geqslant 2^{7}+1$. Hence, we would like to gather some information about the distribution of $M_{2^{7}+1}$. More precisely, we are interested in the behaviour of $\theta_{i}={p_{129,i}}/{p_{129,i+1}}$ for $i\leqslant 11$, which we present in Table 1.
@@
We briefly see a superexponential trend of proportions $\theta_{i}$, so the possibility of using Claim 2 for $n\geqslant 2^{7}+1$ is justified. It might seem that the choice of $n$ is arbitrary, but it occurs that the distribution of $M_{2^{6}+1}$ does not satisfy the necessary assumptions for privacy loss, although $M_{2^{6}+1}$ can still fulfil the property of $(\varepsilon(n),\delta)$-DP with the parameters given in Theorem 1.
@@
Let $k\in\mathbb{N}\setminus\{0\}$ and $n_{k}=2^{k}+1$. If $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$ for every $c$ in the interval [k:4][-k:4], then
@@
Realise that for c=kc=-k, the required inequality is trivial. Therefore, we can safely consider only $c\in[-k+1:4]$. We would like to prove this inductively with respect to $c$ and $N$. Assume that for some $N\geqslant n_{k}$ and any $d\in\{c-1,c\}$ we have $p_{N,k+d}<2^{d+3}p_{N,k+d+1}$.
Then also
@@
[…]
@@
According to the previous discussion on formula (10), we examine
@@
Realise that if $n=2^{l-1}+1$ for some $l\in\mathbb{N}$, then a little adjustment is necessary. Indeed, let now $c-1=k-l\in[-4:4]$, and once again, Lemma 3 provides $p_{n-1,k-1}<~2^{c+2}p_{n-1,k}$. However, it still holds that:
@@
[…]
Theorem 6 only provides $\varepsilon(n)\leqslant-\ln\left(1-{16}/{n}\right)$ for $n>128$ (compare with (10)). However, in Figure 5 we may briefly see that the above inequality is true for smaller numbers of requests $n$ as well.
@@
[…]
Figure 5: Exact values of $\varepsilon(n)$ parameter for $n\leqslant 160$ compared with plots of sequences $\ln(1-{16}/{n})$ and $\ln(1-{8}/{n})$.
In the Figure 5 one can see that values of $\varepsilon(n)$ are strictly between sequences $\ln(1-~{8}/{n})$ and $\ln(1-~{16}/{n})$ for $n\in[17:160]$. We can also observe that $\varepsilon(n)~\approx~2^{4-\lceil\log(n)\rceil}$ in this interval. Note that $\lceil\log(n)\rceil\leqslant 4$ for $n\leqslant 16$, so $\lceil\log(n)\rceil-4<1$, but $M$ is always positive. This can justify the chaotic behaviour of the process for $n\leqslant 16$. Nevertheless, Figure 5 affirms the quality of $\varepsilon(n)$ parameter established in Theorem 1.
Moreover, let us mention that for the strategy used in the proof, we cannot pick smaller $\varepsilon$ of the same form:
@@
As can be seen in Figure 5, the minimal $c$ for $\varepsilon(n)$, where $n$ is of the form $2^{k}$, for $k>5$, is also very close to $16$.
@@ 8.2 General result on Morris’ Counter privacy (Theorem 2)
For our convenience, let us denote $\rho:=\lceil c\log(\ln(n))\rceil$. We assume that $\rho\geqslant 1$. First, we show that $\delta_{1}^{\ast}:=\mathbb{P}\left(M_{n}\leqslant\lceil\log(n)\rceil-\rho-1\right)=O\left(n^{-\left(\ln(n)\right)^{c-1}}\right)$. The proof is analogous to the one of Lemma 1 (we omit similar parts). Indeed,
@@
Now, we are going to prove $\delta_{2}^{\ast}:=\mathbb{P}\left(M_{n}\geqslant\lceil\lg(n)\rceil+\rho+1\right)=O\left(n^{-1}\left(\ln(n)\right)^{-c}\right)$. We use a similar notation and technique as in the proof of Lemma 2, but this time we utilise a discrete discrete two-dimensional simplex $(\rho-1)$:
@@
We couple $(M_{n})_{n}$ with the same process $(X_{n})_{n}$ as in Lemma 2. Roughly speaking, $X_{0}=1$ and $X_{n}$ almost always increments by $1$ until $n=\lceil\log(n)\rceil+1$ and it further follows the same incrementation rule as the Morris counter. Then
@@
In addition, we would like to consider fractions $\frac{p_{n+1,k}}{p_{n,k}}$ for $k\in J_{n}(c)$ as in the proof of Theorem 6. Indeed
@@
Let us denote the above sum by $\varsigma_{k}(n-k+1)$. We note that
@@
Realise that $\varsigma_{k-1}(n-k+2)\leqslant\varsigma_{k-1}(n-k+1)\sum_{i=1}^{k-1}\left(1-2^{-i}\right)=\varsigma_{k-1}(n-k+1)\left(k-2+2^{k-1}\right)$. This follows from the fact that each summand of $\varsigma_{k-1}(n-k+2)$ can be obtained from some summands of $\varsigma_{k-1}(n-k+1)$ by multiplication by one of the terms $\left(1-2^{-i}\right)$. Moreover, note that $\varsigma_{k}(n-k+1)$ has $\binom{n-k+1+(k-1)}{k-1}=\binom{n}{k-1}$ summands and, similarly, $\varsigma_{k-1}(n-k+1)$ has $\binom{n-1}{k-2}$ summands. One can briefly see that a function $f(i)=\left(1-2^{-i}\right)$ is increasing, hence
@@
Due to the monotonicity of $f$, one can use cascading substitutions: some of $f(k)$ by f(k1)f(k-1), then some of f(k1)f(k-1) by f(k2)f(k-2) etc., in order to balance the numbers of all the occurring summands, what provides:
@@ 8.3 Main result for MaxGeo Counter (Theorem 3)
We have $n$ increment requests, which influence the value of MaxGeo Counter $M$. Then the result of the mechanism is $X=\max(X_{1},\ldots,X_{n})$, where $X_{i}\sim\mathrm{Geo}({1}/{2})$ are pairwise independent. First, we observe that if $n=m$, the counter trivially satisfies differential privacy, as the probability distribution of $X$ does not change. From now on, we assume that |nm|=1|n-m|=1. See that
@@
It is easy to see that the above decreases as $n$ increases. Then
@@
where the approximation is the result of the substitution of $l_{\varepsilon}$ without ceiling. $\qed$
@@ 9 Conclusions and Future Work
In this paper, we have investigated probabilistic counters from a privacy-protection perspective. We have shown that Morris Counter and MaxGeo Counter inherently guarantee differential privacy from the mechanism itself, provided that there is at least a small fixed number of increment requests. Otherwise, the counter has too low a value and, intuitively, the result is not randomised enough. We have also shown that the constant in our Morris Counter result cannot be improved further.
@@
We have shown how to perform data aggregation, namely a distributed survey, in a privacy-preserving manner using probabilistic counters. We clarified that this type of solution is especially efficient when one cares about memory resources, like in many Big Data related problems. Note that the security model in this paper was somewhat optimistic. Unfortunately, in such a setting, there is little incentive to use them other than when we already have them deployed and working as aggregators due to e.g., memory-efficiency requirements. However, this would change tremendously if we weakened these assumptions. This seems a very promising way to continue our research from this paper. Namely, we focused on privacy and can still not weaken the security assumptions and allow the Adversary to extract information from channels between users and the aggregator. That would put us in the so-called Local Model, where each user is responsible for the data randomisation. However, such an approach requires us to be able to perform probabilistic counter in an oblivious manner, which, to the best of our knowledge, was not explored before.
@@
In Subsection 4.1, we have mentioned the generalisation of the Morris Counter (for bases $a>1$). Analysis of privacy properties of such variants of Morris Counters and various probabilistic counters presented, for example, in [10], [26] may also be promising directions of further research.
@@
In this paper, we focus on the standard definition of differential privacy. However, there is also an issue of preservation of differential privacy for requests given by a group of $k$ users or one individual sending up to $k$ dependent requests over time. This can be described in the language of the so-called $(\varepsilon,\delta)-k$-DP (see the "group privacy scenario" in [18]) . A group of people may tend to behave in the same manner, so they may send $k$ requests in a row. Especially this ”group” may be represented by a single person colluding with the Adversary. It is worth mentioning that this type of generalisation creates an opportunity to modify probabilistic counters so that each incrementation request executes the update request multiple times to reduce the variance of the rescaled estimator. Intuitively, this extension should be especially efficient in preserving the standard differential privacy property when $\varepsilon(n)=\frac{c}{n}+o(n^{-1})$ (as a parameter of standard differential privacy), because both $c$ and $n$ should scale with $k$ linearly. Hence, the next challenging problem is showing that Morris and MaxGeo Counters satisfy $k$-DP property with similar privacy parameters.
@@
The Morris Counter and the MaxGeo Counter are considered the most popular probabilistic counters. However, the results of this paper shed new light on the properties of the probabilistic counter, in general. There is a possibility to provide analogous differential privacy properties for other probabilistic counters. Moreover, this paper enables the provision of differentially private algorithms for other applications, especially those based on Morris or MaxGeo Counter. For example, in Section 4 we mentioned PCSA and HyperLogLog Counters together with their variances, which can be manually adjusted to the applications. The proper choice of the $m$ parameter implies an exchange of memory usage to improve the accuracy of the estimation. We have mentioned that these counters’ differential privacy parameters can be obtained via Observation 1. However, such a direct result may not be satisfying. Hence a more precise calculation is needed. For instance, Observation 1 may be utilised again with some concentration inequalities.
@@ Acknowledgement
The authors sincerely thank the reviewers and editors for their significant time and effort invested in editing and improving this work. This work was supported by the Polish National Science Center grant number UMO-2018/29/B/ST6/02969.
@@ Appendix A Technical Lemmas and Proofs Related to Differential Privacy of Morris Counter
For the sake of completeness, we present here proofs of all technical lemmas that are not directly connected to Theorem 1. Some of computations are supported by Wolfram Mathematica ver.11.3 ([29]). Whenever we obtain a result in this manner, we indicate it by $\stackrel{{\scriptstyle W}}{{=}}$ sign. Usually results are precise, however in some cases, final forms are attained numerically.
We often struggle with expressions of a pattern 11/y1-{1}/{y}, so let us denote this function as $a(y)$ to abbreviate formulas.
@@
and in a consequence $a(2^{-s}x)^{2x+1}<\exp(-2^{s+1})$ for any reasonable $s$.
Moreover, let us introduce
@@
Let $d:=1-{2^{s}}/{x}$ and realize that $d\in[{3}/{4},1)$ and $2^{s}-1+d\left(-{2^{s}}/{3}-2^{-6}-2^{2s-2}\right)>0$. Indeed, if we put $z=2^{s}$, then we attain a quadratic inequality in $z$ variable, with determinant $\Delta=1-\frac{5d}{3}+\frac{55d^{2}}{576}$, that is negative for $d\in[{3}/{4},1)$.
@@
[…]
where $\mathrm{cond}]\!]$ is the Iverson bracket of the condition $\mathrm{cond}$.
Realize that for $t\in\{0,1\}$ and $k\geqslant 5$, $|\tau(k,t)+\kappa(k,t)|<2^{-50}<10^{-15}$. Now, consider the differences between the consecutive elements of sequences:
@@
$=$ $\hskip-8.53581pt\sum_{i=0}^{\left\lfloor\frac{k+t+2}{2}\right\rfloor}\hskip-5.69054pt2^{-i(2i-1)}r_{2i+1}r_{k+t+4-2i}\left[a(2^{2i+1})\left(a(2^{5+t-2i}x)^{2x+1}-a(2^{4+t-2i}x)^{x+2}\right)\right$.
@@
$\stackrel{{\scriptstyle W}}{{=}}a(2^{2i+1})a(u_{t}x)^{2x+1}\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{u_{t}+1}{u_{t}^{3}x}-\frac{1}{u_{t}^{3}x^{2}}\right)$
@@
$\stackrel{{\scriptstyle W}}{{=}}a(2^{2i+1})\exp\left(-\frac{2}{u_{t}}\right)\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{u_{t}+1}{xu_{t}^{3}}-\frac{1}{x^{2}u_{t}^{3}}\right)$ (12)
@@
$\stackrel{{\scriptstyle W}}{{=}}$ $a(2^{2i+1})a\left(\frac{u_{t}}{2}x\right)^{x-1}\frac{1}{x}\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{10u_{t}+1}{u_{t}^{3}x}+\frac{8}{u_{t}^{3}x^{2}}\right)$ (13)
@@
$\stackrel{{\scriptstyle W}}{{=}}\frac{a(2^{2i+1})}{x}\exp\left(-\frac{2}{u_{t}}\right)\left(\frac{3u_{t}+1}{u_{t}^{2}}-\frac{2+5u_{t}+4u_{t}^{2}}{u_{t}^{4}x}-\frac{2+4u_{t}-\frac{575}{32}u_{t}^{2}+\frac{387}{32}u_{t}^{3}}{u_{t}^{6}x^{2}}\right$.
@@
[…]
@@
$\sum_{i=5}^{\left\lfloor\frac{k+2}{2}\right\rfloor}2^{-i(2i-1)}r_{2i+1}a(2^{5-2i}x)^{2x+1}\leqslant\frac{R\;2^{-45}\exp(-64)}{1-2^{-21}\exp(-192)}=1.5784\ldots\cdot 10^{-41}~,$
@@
However, according to Theorem 4, we also present the numerical values of the sequence $(p_{2^{k}+1,k+4})_{k=2}^{14}$ in the Table 2. We can now easily see that for any $k\geqslant 2$ we attained $p_{2^{k+1}+1,k+5}-p_{2^{k}+1,k+4}<0$ .